Cost of Equity with and without Flotation

Jarett & Sons's common stock currently trades at $26.00 a share. It is expected to pay an annual dividend of $1.75 a share at the end of the year (D1 = $1.75), and the constant growth rate is 5% a year.

  1. What is the company's cost of common equity if all of its equity comes from retained earnings? Round your answer to two decimal places. Do not round your intermediate calculations.
    %
  2. If the company issued new stock, it would incur a 11% flotation cost. What would be the cost of equity from new stock? Round your answer to two decimal places. Do not round your intermediate calculations.
Answer #1

a.Cost of equity=(D1/Current price)+Growth rate

=(1.75/26)+0.05

=11.73%(Approx).

b.Cost of equity=(D1/Current price(1-Flotation cost)+Growth rate

=(1.75/(26*(1-0.11))+0.05

=(1.75/23.14)+0.05

=12.56%(Approx).